Two charged spheres, q1= -2.00 x 10 -5 C and q2 = -1.50 x 10 -5 C are 0.1 m apart. Calculate the force that each charge exerts on the third charge, q3 = 5.00 x 10 -6 C which is 0.06 m away from q2.
1
Expert's answer
2020-12-14T07:16:10-0500
"F=k\frac{q_1q_3}{d-x}-k\frac{q_2q_3}{x}\\
F=9\cdot10^9\frac{(2\cdot10^{-5})(5\cdot10^{-5})}{0.1-0.06}-9\cdot10^9\frac{(1.5\cdot10^{-5})(5\cdot10^{-5})}{0.06}\\F=112.5\ N"